# Imbalance in Balance: Online Concept Balancing in Generation Models

### v1

@@ 5.1 Quantitative Comparison
As shown in Table 3, we compare baseline, A&E [5], finetuning, and from-scratch training of our IMBA loss on LC-Mis [60], T2i-CompBench [18] and ours Inert-CompBench. Compared with the diffusion loss in the baseline, our IMBA loss can significantly improve the concept composition ability, whether training from scratch or finetuning from a pre-trained model. In addition, training from scratch yields better results compared to fine-tuning. Besides, when our approach achieves similar improvements in object missing, A&E [5] performs slightly better on CLIP score, but far worse on attribute leakage (shape,color,texture,VQA). This is because A&E is limited by the foundational generation model and cannot generate concepts that the generation model does not understand. Meanwhile, since the improvement of fine-tuning on Inert-CompBench is limited, it indicates that inert concepts require a longer training process to enhance concept composition ability.
@@ B.2 Comparison with Frequecy-based Method
Since the text is a joint distribution of multiple concepts, it is difficult to calculate weights from a frequency perspective, and there is little concept-balancing work for text-to-image generation. Therefore, we compare IMBA loss with a frequency based method on class-image [9]. We sample 5 concepts each from the head and tail concepts and combine the data containing these concepts in the training set into a new subset. We then finetune the model on the subset using the frequency-based and our method respectively. Meanwhile, we pair the 10 concepts to generate 5 captions for each pair as the test set. As shown in the Table 4, our method outperforms the frequency-based method.

### v2

@@ 5.1 Quantitative Comparison
As shown in Table 3, we compare baseline, A&E [chefer2023attend], finetuning, and from-scratch training of our IMBA loss on LC-Mis [zhao2024lost], T2i-CompBench [huang2023t2icompbench] and ours Inert-CompBench. Compared with the diffusion loss in the baseline, our IMBA loss can significantly improve the concept composition ability, whether training from scratch or finetuning from a pre-trained model. In addition, training from scratch yields better results compared to fine-tuning. Besides, when our approach achieves similar improvements in object missing, A&E [chefer2023attend] performs slightly better on CLIP score, but far worse on attribute leakage (shape,color,texture,VQA). This is because A&E is limited by the foundational generation model and cannot generate concepts that the generation model does not understand. Meanwhile, since the improvement of fine-tuning on Inert-CompBench is limited, it indicates that inert concepts require a longer training process to enhance concept composition ability.
@@ B.2 Comparison with Frequecy-based Method
Since the text is a joint distribution of multiple concepts, it is difficult to calculate weights from a frequency perspective, and there is little concept-balancing work for text-to-image generation. Therefore, we compare IMBA loss with a frequency based method on class-image [cui2019classloss]. We sample 5 concepts each from the head and tail concepts and combine the data containing these concepts in the training set into a new subset. We then finetune the model on the subset using the frequency-based and our method respectively. Meanwhile, we pair the 10 concepts to generate 5 captions for each pair as the test set. As shown in the Table 4, our method outperforms the frequency-based method.